Use the common-size financial statements found here:

6

parts remaining

Clear All

Check Answer

Data Table

Common-Size Balance Sheet

2016

Cash and marketable securities

$

500

1.5

%

Accounts receivable

6,000

18.2

Inventory

9,500

28.8

Total current assets

$

16,000

48.5

%

Net property, plant, and equipment

17,000

51.5

Total assets

$

33,000

100.0

%

Accounts payable

$

7,200

21.8

%

Short-term notes

6,800

20.6

Total current liabilities

$

14,000

42.4

%

Long-term liabilities

7,000

21.2

Total liabilities

$

21,000

63.6

%

Total common shareholders’ equity

12,000

36.4

Total liabilities and shareholders’ equity

$

33,000

100.0

%

Common-Size Income Statement

2016

Revenues

$

30,000

100.0

%

Cost of goods sold

(20,000)

66.7

Gross profit

$

10,000

33.3

%

Operating expenses

(8,000)

26.7

Net operating income

$

2,000

6.7

%

Interest expense

(900)

3.0

Earnings before taxes

$

1,100

3.7

%

Income taxes

(400)

1.3

Net income

$

700

2.3

%

LOADING...

to respond to your boss' request that you write up your assessment of the firm's financial condition. Specifically, write up a brief narrative that responds to the following questions:

a. How much cash does Patterson have on hand relative to its total assets?

b. What proportion of Patterson's assets has the firm financed using short-term debt? Long-term debt?

c. What percent of Patterson's revenues does the firm have left over after paying all of its expenses (including taxes)?

d. Describe the relative importance of Patterson's major expense categories, including cost of goods sold, operating expenses, and interest expenses.

a. How much cash does Patterson have on hand relative to its total assets?

The cash Patterson has on hand relative to its total assets is

nothing %.

(Round to one decimal place.)
Answer #1

The common size statements found is correct as per question given
a)=500/33000=1.5%
b)short term debt=6800/33000=20.6%
long term debt=7000/33000=21.2%
c)Here there are asking net income % and it is =(700/30000)=2.3%
d)Cost of good sold is 66.7% of revenues, operating expenses is 26.7% of revenues and interest expense is just 3% of revenues